OpenStudy (jadedry):

Hi! I'm having trouble with the following question: "Solve the following quadratic formula by completing the square: x^2 -2x +a = 0 " So far, I'm managed to find out that a)The equation has 2 distinct roots because "b^2 -4ac" = positive b) One of the roots is 1. My textbook says that the other answer is sqrt(1-a) but I can't get that answer. My calculations show that the answer should be (sqrt(1-a) + 1 ) Could you please show working? Thanks in advance!

hartnn (hartnn):

so you know how to solve quadratics using completing the square method... can you just post first few steps of your working??

hartnn (hartnn):

i believe you reached till \(\large x^2 -2x+1 = 1-a\) right ??

hartnn (hartnn):

oh and just so you know, your answer of \(\large x = \sqrt{1-a}+ 1 \) is correct. and the book is wrong

hartnn (hartnn):

another answer is \(\large x = -\sqrt{1-a} +1 \)

OpenStudy (jadedry):

I did indeed reach \[x^{2}-2x +1 = 1-a\] at which point it seemed that \[x-1 = \sqrt{1-a}\] -- Glad to know that my answer was right though! I was ripping my hair out. Thanks for clarifying. c:

hartnn (hartnn):

welcome ^_^ i see you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (jadedry):

Haha, thank you for warm welcome! I'll probably be lurking around in these forums quite often. c: